### v3

@@ 6.3 Cryptocurrency Trading Tasks
Table 7: Cryptocurrency trading task performance. The second-level LOB data for Bitcoin is split into out-of-sample data for training and in-sample data for testing. Ensemble models use majority voting on agent actions.
@@
The FinRL Contest 2024 Cryptocurrency Trading with Ensemble Learning and the FinRL Contest 2025 FinRL-AlphaSeek for Crypto Trading are tasks designed for Bitcoin trading.
6.3.1 Cryptocurrency Trading with Ensemble Methods
@@
In our experiment, the cryptocurrency trading task for Bitcoin (BTC) $(K=1)$ is performed using three ensemble models, and individual DQN, Double DQN, and Dueling DQN agents.
@@
Cryptocurrency data: The dataset comprises second-level LOB data for BTC from 04/07/2021 11:32:42 to 04/19/2021 09:54:22. Adaptations from 101 formulaic alphas 30 are calculated based on LOB data to extract insights into market behaviors, such as momentum, mean-reversion, and market anomalies. A recurrent neural network (RNN) further processes the 101 alphas into $I=8$ technical indicators. It reduces the complexity of the input data and enhances the ability to predict market trends, thus improving generalization and avoiding overfitting. The RNN is trained on data from 04/07/2021 11:32:42 to 04/17/2021 00:38:02 without future information leaks. The agents are trained on the in-sample data from 04/17 00:38:03 to 04/19 09:09:21 and tested on the out-of-sample data from 04/19 09:09:22 to 04/19 09:54:22.
@@
Agents for cryptocurrency trading task. We use DQN, Double DQN, and Dueling DQN agents. The policy network for each agent consists of a feed-forward neural network with three 128-unit hidden layers. We set an exploration rate of $0.005$, a learning rate of $2\cdot 10^{-6}$, and a batch size of $512$.
@@
Ensemble methods The first method (Ensemble 1) consists of 11 DQN, $1$ Double DQN, and $1$ Dueling DQN agents; the second method (Ensemble 2) consists of $3$ DQN, $3$ Double DQN, and $3$ Dueling DQN agents; the third method (Ensemble 3) consists of $10$ agents for each type. As in Section 5.1.2, three ensemble models use a majority voting approach to aggregate the agents’ actions. All ensemble models and individual agents are trained on the in-sample data and tested on the out-of-sample data.
@@
Results. As seen in Table 7 and Fig. 7 (b), Double DQN and Dueling DQN agents have similar performance, with cumulative returns of $0.48\%$. This is lower than the BTC price baseline. Despite this, they achieve higher Sharpe ratios of $0.21$ and lower maximum drawdowns of $0.98\%$ than the fixed-time exit strategy and BTC price baseline, suggesting effective risk management. The three different ensemble models have similar performances, which may be due to the limited action space at each timestep, causing agents to output identical actions. Their cumulative returns are close to the BTC price baseline. Moreover, the ensemble models outperform all individual agents in all metrics, achieving the highest Sharpe ratio of $0.28$ and the lowest maximum drawdown of $0.73\%$. They also have the highest win/loss ratio of $1.62$. This shows that ensemble methods can mitigate the risks associated with the decision-making failures of single agents.
@@
We observe that individual agents and majority voting ensembles have near identical performances; in the restricted action space, agent policies may be converging indicating a greater risk of significant losses due to less diversified trading actions. Compared to the spot price and fixed-time exit, the ensemble consistently achieved a higher return over the maximum drawdown ratio, highlighting superior risk-adjusted returns. Due to a lack of diversity in agents, we observed near-identical results between ensembles and individual agents. Nonetheless, over a relatively short timespan of $30$ minutes, we observe that the vectorized agents can outperform other strategies.
@@
This was crucial in maintaining portfolio stability amid high market volatility. The ensemble generally showed a higher win rate and a lower loss rate than the individual agents, reflecting its enhanced decision-making accuracy and consistency. The comparative analysis between the ensemble model and individual trading agents underscores the ensemble’s superior capability in managing risks and capitalizing on market opportunities. While individual agents provide valuable insights and are crucial components of the ensemble, the aggregated approach of the ensemble offers a more robust and effective solution for trading in the volatile cryptocurrency markets.
6.3.2 Cryptocurrency Trading in FinRL Contests
@@
Table 8: The performance of winners for FinRL Contest 2024 Task 1 Cryptocurrency Trading with Ensemble Learnig and FinRL Contest 2025 Task 2 FinRL-AlphaSeek for Crypto Trading.
@@
  • Mt.Everest added new technique indicators in the state of the market environment, including order book imbalance, spread-based liquidity, order flow imbalance, price impact of trades, market-to-limit order ratio, intraday volatility, momentum scores, and noise-to-signal ratio. In addition, the team used principal component analysis (PCA) to extract the top 10 strongest factors from 101 alpha factors 30. To construct the agent pool, they applied different policy networks, including LSTM, transformer, and dilated CNNs. The action was determined through majority voting.
@@
In addition, FinRL-Cryptohttps://github.com/AI4Finance-Foundation/FinRL_Crypto 31 introduces a multi-cryptocurrency trading strategy. The agents were trained using five-minute-level OHLCV data for $K=10$ cryptocurrencies from 02/02/2022 to 04/30/2022. The agents were trained with multiple hyperparameter settings across 10 training-validation splits. To address backtest overfitting, agents with a high probability of overfitting were rejected at a $10\%$ significance level. The model was evaluated on the dataset from 05/01/2022 to 06/27/2022, which includes two market crashes. The best-performing PPO agent achieved a cumulative return of $34.96\%$, outperforming the S&P Cryptocurrency Broad Digital Market Index (S&P BDM Index, $50.78\%$) and equal-weight strategy ($47.78\%$). It also had lower volatility ($0.0020$), compared with S&P BDM Index ($0.0581$) and equal-weight strategy ($0.0042$).
@@ 7 Organizational Aspects
We held FinRL contests at several academic conferences, including ACM International Conference on AI in Finance (ACM ICAIF 2023, 2024) and IEEE International Conference on Intelligent Data and Security (IEEE IDS 2025). In this section, we summarize the organizational experience of the FinRL Contests, including participation, contest process, platform setup, promotion, registration and submission approaches, and communication.
Participation. From 2023 to 2025, a total of 200+ students, researchers, and practitioners participated. Fig. 8 shows the distribution of participation by geography and affiliation type. Among the 141 participants who reported their institutions, $37\%$ participants are from North America, including the United States and Canada. $46\%$ participants come from Asia, mainly from China, South Korea, and India. Among all 200 participants, $71.5\%$ of them are from 62 academic institutions, such as Columbia University, National University of Singapore, and Shenzhen University. $17.5\%$ of them are from 21 industrial institutions.
@@
  • Group chat. We set up a Discord channel of around 500 people and 2 WeChat groups of 260 people to facilitate participant networking, discussion, and direct Q&A.
@@ 8 Conclusions and Future Works
We thank Vatic Investment and Shanghai Frontiers Science Center of Artificial Intelligence and Deep Learning at NYU Shanghai for providing registration fees for participant teams.
Keyi Wang, Nikolaus Holzer, Jiechao Gao, Anwar Walid and Xiao-Yang Liu Yanglet acknowledge the support from Columbia’s SIRS and STAR Program, as well as The Tang Family Fund for Research Innovations in FinTech, Engineering, and Business Operations. Xiao-Yang Liu Yanglet also acknowledges the support from a NSF IUCRC CRAFT center research grant (CRAFT Grant 22017) for this research. The opinions expressed in this publication do not necessarily represent the views of NSF IUCRC CRAFT.
@@
The authors declare no potential conflict of interests.

### v4

@@ 6.3 Crypto Trading Tasks
Table 7: crypto trading task performance. The second-level LOB data for Bitcoin is split into out-of-sample data for training and in-sample data for testing. Ensemble models use majority voting on agent actions.
@@
The FinRL Contest 2024 crypto Trading with Ensemble Learning and the FinRL Contest 2025 FinRL-AlphaSeek for Crypto Trading are tasks designed for Bitcoin trading.
6.3.1 Crypto Trading with Ensemble Methods
@@
In our experiment, the crypto trading task for Bitcoin (BTC) $(K=1)$ is performed using three ensemble models, and individual DQN, Double DQN, and Dueling DQN agents.
@@
crypto data: The dataset comprises second-level LOB data for BTC from 04/07/2021 11:32:42 to 04/19/2021 09:54:22. Adaptations from 101 formulaic alphas 29 are calculated based on LOB data to extract insights into market behaviors, such as momentum, mean-reversion, and market anomalies. A recurrent neural network (RNN) further processes the 101 alphas into $I=8$ technical indicators. It reduces the complexity of the input data and enhances the ability to predict market trends, thus improving generalization and avoiding overfitting. The RNN is trained on data from 04/07/2021 11:32:42 to 04/17/2021 00:38:02 without future information leaks. The agents are trained on the in-sample data from 04/17 00:38:03 to 04/19 09:09:21 and tested on the out-of-sample data from 04/19 09:09:22 to 04/19 09:54:22.
@@
Agents for crypto trading task. We use DQN, Double DQN, and Dueling DQN agents. The policy network for each agent consists of a feed-forward neural network with three 128-unit hidden layers. We set an exploration rate of $0.005$, a learning rate of $2\cdot 10^{-6}$, and a batch size of $512$.
@@
Ensemble methods The first method (Ensemble 1) consists of $1$ DQN, $1$ Double DQN, and $1$ Dueling DQN agents; the second method (Ensemble 2) consists of $3$ DQN, $3$ Double DQN, and $3$ Dueling DQN agents; and the third method (Ensemble 3) consists of $10$ agents for each type. As in Section 5.1.2, three ensemble models use a majority voting approach to aggregate the agents’ actions. All ensemble models and individual agents are trained on the in-sample data and tested on the out-of-sample data.
@@
Results. As seen in Table 7 and Fig. 7 (b), Double DQN and Dueling DQN agents have similar performance, with cumulative returns of $0.48\%$. This is lower than the BTC price baseline. Despite this, they achieve higher Sharpe ratios of $0.21$ and lower maximum drawdowns of $-0.98\%$ than the fixed-time exit strategy and BTC price baseline, suggesting effective risk management. The three different ensemble models have similar performances, which may be due to the limited action space at each timestep, causing agents to output identical actions. Their cumulative returns are close to the BTC price baseline. Moreover, the ensemble models outperform all individual agents in all metrics, achieving the highest Sharpe ratio of $0.28$ and the lowest maximum drawdown of $-0.73\%$. They also have the highest win/loss ratio of $1.62$. This shows that ensemble methods can mitigate the risks associated with the decision-making failures of single agents.
@@
We observe that individual agents and majority voting ensembles have near identical performances; in the restricted action space, agent policies may be converging, indicating a greater risk of significant losses due to less diversified trading actions. Compared to the spot price and fixed-time exit, the ensemble consistently achieved a higher return over the maximum drawdown ratio, highlighting superior risk-adjusted returns. Due to a lack of diversity in agents, we observed near-identical results between ensembles and individual agents. Nonetheless, over a relatively short timespan of $30$ minutes, we observe that the vectorized agents can outperform other strategies.
@@
This was crucial in maintaining portfolio stability amid high market volatility. The ensemble generally showed a higher win rate and a lower loss rate than the individual agents, reflecting its enhanced decision-making accuracy and consistency. The comparative analysis between the ensemble model and individual trading agents underscores the ensemble’s superior capability in managing risks and capitalizing on market opportunities. While individual agents provide valuable insights and are crucial components of the ensemble, the aggregated approach of the ensemble offers a more robust and effective solution for trading in the volatile crypto markets.
6.3.2 Crypto Trading in FinRL Contests
@@
Table 8: The performance of winners for FinRL Contest 2024 Task 1 “crypto Trading with Ensemble Learning” and FinRL Contest 2025 Task 2 “FinRL-AlphaSeek for Crypto Trading”.
@@
  • Mt.Everest added new technique indicators in the state of the market environment, including order book imbalance, spread-based liquidity, order flow imbalance, price impact of trades, market-to-limit order ratio, intraday volatility, momentum scores, and noise-to-signal ratio. In addition, the team used principal component analysis (PCA) to extract the top 10 strongest factors from 101 alpha factors 29. To construct the agent pool, they applied different policy networks, including LSTM, transformer, and dilated CNNs. The action was determined through majority voting.
@@
In addition, FinRL-Cryptohttps://github.com/AI4Finance-Foundation/FinRL_Crypto 30 introduces a multi-crypto trading strategy. The agents were trained using five-minute-level OHLCV data for $K=10$ cryptocurrencies from 02/02/2022 to 04/30/2022. The agents were trained with multiple hyperparameter settings across 10 training-validation splits. To address backtest overfitting, agents with a high probability of overfitting were rejected at a $10\%$ significance level. The model was evaluated on the dataset from 05/01/2022 to 06/27/2022, which includes two market crashes. The best-performing PPO agent achieved a cumulative return of $-34.96\%$, outperforming the S&P crypto Broad Digital Market Index (S&P BDM Index, $-50.78\%$) and the equal-weight strategy ($-47.78\%$). It also had lower volatility ($0.0020$), compared with S&P BDM Index ($0.0581$) and the equal-weight strategy ($0.0042$).
@@ 7 Organizational Aspects
We held FinRL contests at several academic conferences, including ACM International Conference on AI in Finance (ACM ICAIF 2023, 2024), IEEE International Conference on Intelligent Data and Security (IEEE IDS 2025), and IEEE International Conference on Cyber Security and Cloud Computing (IEEE CSCloud 2025). In this section, we summarize the organizational experience of the FinRL Contests, including encouraging participation, contest process, platform setup, event promotion, registration and submission approaches, and communication.
Participation. From 2023 to 2025, a total of 200+ students, researchers, and practitioners participated. Fig. 8 shows the distribution of participation by geography and type of affiliation. Among the 141 participants who reported their institutions, $37\%$ of participants are from North America, including the United States and Canada. $46\%$ of participants come from Asia, mainly from China, South Korea, and India. Among the 200+ participants, $71.5\%$ of them come from 62 academic institutions, such as Columbia University, the National University of Singapore, and Shenzhen University. $17.5\%$ of them are from 21 industrial institutions. (The statistics exclude the ongoing FinAI Contest 2025 @ IEEE CSCloud).
@@
  • Group chat. We set up a Discord channel of 500+ people and 2 WeChat groups of 270+ people to facilitate participant networking, discussion, and direct Q&A.
@@ 8 Conclusions and Future Works
The authors thank Vatic Investment and the Shanghai Frontiers Science Center of Artificial Intelligence and Deep Learning at NYU Shanghai for covering registration fees for participant teams.
Keyi Wang, Nikolaus Holzer, Jiechao Gao, Anwar Walid, and Xiao-Yang Liu Yanglet acknowledge the support from Columbia’s SIRS and STAR Program, as well as The Tang Family Fund for Research Innovations in FinTech, Engineering, and Business Operations. Xiao-Yang Liu Yanglet also acknowledges the support from a NSF IUCRC CRAFT center research grant (CRAFT Grant 22017) for this research. The opinions expressed in this publication do not necessarily represent the views of NSF IUCRC CRAFT.
@@
The authors declare no potential conflict of interest.